Two cards are drawn with replacement. $X$ = number of aces. The probability $P[X \geq 1]$ equals:
A$25/169$
B$1/169$
C$144/169$
D$24/169$
Answer & Solution
Correct answer: A. $25/169$
$P[X \geq 1] = 1 - P[X=0] = 1 - 144/169 = 25/169$.
(Or compute directly: $P[X=1] + P[X=2] = 24/169 + 1/169 = 25/169$.)
Keep practising Arithmetic Progression
Browse all Arithmetic Progression questions for GMAT, or explore more Mathematics topics.
Related questions